# Leveraging Content and Acoustic Representations for Efficient Speech Emotion Recognition

### v1

@@ III-B2 Loss
An utterance level loss, $L_{utt}$ and a frame-level acoustic loss, $L_{acoust}$, are employed for training the semantic and acoustic arms, respectively. Denoting the mean-pooled output from the pre-trained RoBERTa model by $\boldsymbol{y}_{text}$ and the output from the semantic arm as $\hat{\boldsymbol{y}}_{sem}$, the semantic loss is the mean square error (MSE) loss:
@@
For the frame level loss, let $\hat{\boldsymbol{y}}_{acoust}\in R^{N\times T\times D}$ denote the output of the acoustic arm, where $N$, $T$ and $D$ denote the batch size, number of frames per utterance and the dimension of the representation, respectively. The loss is defined as:
$L_{acoust}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}||\boldsymbol{y}_{pase}^{ij}-\hat{\boldsymbol{y}}_{acoust}^{ij}||_{2}^{2}$ (2)
where $\boldsymbol{y}_{pase}$ denotes the acoustic target from the PASE model.
@@ III-B3 Inference
For the different downstream tasks, the model is evaluated according to the paradigm proposed in SUPERB benchmark [49]. The representations from the semantic and acoustic arms are concatenated layer-wise. A convex combination of the layer representations is fed to a classification head. The only learnable parameters on the downstream task are the convex combination weights as well as the light-weight classification head, while the weights of the common, acoustic and semantic arms are kept frozen for the downstream supervised training task. This setting is also followed for other baseline models compared in this work.
@@ IV-A Pre-training
The MSP-PODCAST corpus [50] is used for the task of pre-training. A total of $149,307$ speech-turns amounting to $230$ hours of emotional speech data is used. Out of the total number of samples, $80\%$ of the data is randomly chosen as the training set while the remaining $20\%$ serves as the validation set. As the pre-training in the proposed CARE model is unsupervised, only the raw audio data is required. The Whisper-large-v3 model [^1] https://huggingface.co/openai/whisper-large-v3 is used for generating the transcripts, while the pre-trained RoBERTa model [^2] https://huggingface.co/FacebookAI/roberta-base is used for encoding the text data. The common arm is initialized with first $6$ layers of the WavLM base model [^3] https://huggingface.co/microsoft/wavlm-base, while the acoustic arm is initialized with the last $6$ layers of the WavLM base model.
TABLE I: Summary of the evaluation datasets. Class balance denotes if the datasets are balanced across classes. The last column indicates if the training/test data have common speakers. For all the datasets, we perform $5$-fold evaluation.
@@
TABLE II: Results on the different downstream datasets in terms of weighted F1-score. All numbers are averaged over $5$ random initializations of the downstream network. We also show the results of the different components of CARE in this table.
@@
Semantic Arm Acoustic Arm CARE
@@
IV-B Downstream tasks
IV-B1 IEMOCAP
The IEMOCAP dataset consists of $151$ video recordings split into 5 sessions. Each of these sessions is a conversation between a pair of subjects, one male and one female. Each recording is split into multiple utterances. There are a total of $10,039$ utterances, each of which is labeled by human annotators as belonging to one of the $10$ emotions - “angry”, “happy”, “sad”, “neutral”, “frustrated”, “excited”, “fearful”, “surprised”, “disgusted” or “other”. Keeping in line with previous works, we do a four-way classification task where we consider “angry”, “happy”, “sad”, “neutral” and “excited” categories (with “excited” and “happy” categories merged). We also have a separate setting of $6$ emotional classes [59]. The first $6$ emotion classes are considered for this setting.
IV-B2 MELD
The MELD dataset [52] is a multi-party dataset created from video clippings of the popular TV show, “Friends”. A seven way classification task is performed on this dataset, with each utterance being labeled as one of the $7$ emotions - “angry”, “sad”, “joy”, “neutral”, “fear”, “surprise” or “disgust”.
IV-B3 CMU-MOSI
The CMU-MOSI dataset [53] has a total of $2199$ utterances. Each utterance is labeled in the range of $[-3,3]$. Following previous works, we treat this as a binary classification problem with utterances having sentiment values in the range $[-3,0)$ being classified as negative sentiment and those with values in the range $[0,3]$ considered as positive sentiment. The dataset partitioning follows the prior work by Poria et. al [60].
IV-B4 DAIC-WOZ
The DAIC-WOZ dataset [54] is a benchmark dataset for depression detection, consisting of $189$ clinical interviews between the patient and the interviewer. Out of these $189$ interviews, $107$ are part of the training set while $35$ interviews are part of the development subset. The dataset suffers from a data imbalance problem, with only $30$ interviews labeled as “depressed” in the train set . In order to increase the balance, we follow [61] and extract $100$ utterances randomly from each interview, which is labeled as depressed, while only $39$ utterances are selected for interviews classified as “normal”. The utterances from each interview are chosen randomly for the $5$ splits. Following prior work [62, 63, 64, 65, 66, 61], we report results on the development set of this dataset.
IV-B5 RAVDESS-Song
The RAVDESS-Song dataset [55] has a total of $1012$ song recordings sung by $23$ different singers. Each recording in this dataset is sung in one of six different emotions, namely, “neutral”, “calm”, “happy”, “sad”, “angry” and “fear”. We conduct a speaker independent evaluation for this dataset, and create $5$ different splits. For each split, we keep recordings from $16$ singers for training, while recordings from $3$ separate singers are used for validation. The recordings from the remaining $4$ speakers are used for evaluation of the models.
IV-B6 CaFE
IV-B7 EmoDB
The EmoDB dataset [57] has a total of $535$ utterances spoken by 10 different speakers. Each utterance in this dataset is categorized as one of the seven emotions - “neutral”, “angry”, “disgust”, “sad”, “boredom”, “fear” and “joy”. Similar to RAVDESS-Song and CaFE, we create $5$ different speaker independent splits for this dataset. The utterances belonging to $6$ different speakers are chosen for training while the remaining $4$ speakers are used for validation and testing equally. The speakers used for train, validation and testing are chosen randomly for each of the $5$ splits.
IV-C Loss and evaluation metrics
Since these datasets are categorical in natural, we use the cross entropy loss for training the downstream model weights. For testing, we use the weighted F1-score as the evaluation metric. Denoting the F1 score of a class $c$ with $N_{c}$ samples, by $F1_{c}$, the weighted F1-score is defined as:
$WF1=\frac{1}{\sum_{c=1}^{C}N_{c}}\sum_{c=1}^{C}N_{c}\times F1_{c}$ (3)
IV-D Implementation details
Pre-training: During pre-training, all the speech utterances from MSP-PODCAST are padded or randomly cropped to a duration of $5$ seconds. The model is trained with a learning rate of $1e$-$5$ and a batch size of $128$ with AdamW [67] as the optimizer. The model is trained for a total of $200,000$ steps and the best model parameters based on validation set performance are chosen for evaluation of downstream datasets.
Fine-tuning and evaluation: For the downstream task training, the speech signals are cropped to a maximum duration of $30$ seconds or padded to a minimum duration of $1$ second. For the depression detection dataset, DAIC-WOZ, each speech segment has a duration of $10$ seconds [61]. The representations from the layers are accumulated using a convex combination with learnable weights. The classification head is a two-layer feed-forward neural network with the ReLU activation [68]. Only the weights corresponding to convex combination of different layer representations and those in the two-layer fully connected classification head are trained for each downstream dataset, in line with the framework of SUPERB [49]. We use a batch size of $32$ with a learning rate of $1e$-$4$ and train the model for $50$ epochs. The AdamW optimizer is used here as well. All the models, including the CARE and the baseline systems, utilize the same data and classification backend, without any additional hyper-parameter tuning or model selection. Thus, the design allows the fair comparison of the representations.
IV-E Results
The results of our model across various downstream datasets are presented in Table II.

### v2

@@ III-B2 Loss
A semantic loss, $L_{sem}$ and a frame-level acoustic loss, $L_{acoust}$, are employed for training the semantic and acoustic encoders, respectively. Denoting the semantic supervision by $\boldsymbol{y}_{text}$ and the output from the semantic encoder as $\hat{\boldsymbol{y}}_{sem}$, the semantic loss is the mean square error (MSE) loss:
@@
For the frame level loss, let $\hat{\boldsymbol{y}}_{acoust}\in\mathbb{R}^{N\times T\times D}$ denote the output of the acoustic encoder, where $N$, $T$ and $D$ denote the batch size, number of frames per utterance and the dimension of the representation, respectively. The loss is defined as:
$L_{acoust.}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}||\boldsymbol{y}_{pase}^{ij}-\hat{\boldsymbol{y}}_{acoust}^{ij}||_{2}^{2}$ (2)
where $\boldsymbol{y}_{pase}$ denotes the acoustic target.
@@ III-B3 Inference
For evaluating the model across various downstream tasks, we adopt the paradigm proposed in the SUPERB benchmark [41]. The outputs from each transformer layer of the acoustic encoder are concatenated with the outputs from the convolution block following each transformer layer in the semantic encoder. These are then combined with layer-wise outputs from the common encoder and the convolutional feature extractor. This process yields a total of $13$ layer representations—one from the convolutional feature extractor, six from the common encoder, and six from the concatenated semantic and acoustic encoders. A convex combination of these layer representations is then fed into a classification head. It is to be noted that, during inference, the fully-connected layer in the acoustic encoder and the average pooling block in the semantic encoder are not used.
In this setup, the only learnable parameters for the downstream tasks are the weights for the convex combination and those of the lightweight classification head; the weights of the common, acoustic, and semantic encoders remain frozen during supervised training.
@@ IV-A Pre-training
The usage of podcast recordings for pre-training has already been motivated in Sec. III. Raw audio data from the MSP-PODCAST corpus [42] is used for the task of pre-training. A total of $149,307$ speech-turns amounting to $230$ hours of emotional speech data are used. Out of the total number of samples, $80\%$ of the data is randomly chosen as the training set while the remaining $20\%$ serves as the validation set. The Whisper-large-v3 model [^1] https://huggingface.co/openai/whisper-large-v3 is used for generating the transcripts (the WER observed is $12.53\%$), while the pre-trained RoBERTa model [^2] https://huggingface.co/FacebookAI/roberta-base is used for encoding the transcripts. The common encoder is initialized with first $6$ layers of the WavLM-base model [^3] https://huggingface.co/microsoft/wavlm-base, while the acoustic encoder is initialized with the last $6$ layers of the same. The convolutional feature extractor is also initialized from the WavLM-base model. The $6$ transformer layers of the semantic encoder are initialized with the weights of the last $6$ layers of a pre-trained RoBERTa base model, while the convolutional adapters are randomly initialized.
TABLE I: Summary of the evaluation datasets. Class balanced denotes if the datasets are balanced across classes. The last column indicates if the training/test data have common speakers. For all the datasets, we perform $5$-fold evaluation.
@@
IV-B Downstream Tasks
IV-B1 IEMOCAP
The IEMOCAP dataset consists of $151$ video recordings split into 5 sessions. Each of these sessions is a conversation between a pair of subjects. Each recording is split into multiple utterances. There are a total of $10,039$ utterances, each of which is labeled by human annotators as belonging to one of the $10$ emotions - “angry”, “happy”, “sad”, “neutral”, “frustrated”, “excited”, “fearful”, “surprised”, “disgusted” or “other”. Keeping in line with previous works, we do a four-way classification task where we consider “angry”, “happy”, “sad”, “neutral” and “excited” categories (with “excited” and “happy” categories merged). We also have a separate setting of $6$ emotional classes [50]. The first $6$ of the $10$ emotion classes are considered for this setting.
IV-B2 MELD
The MELD dataset [44] is a dataset created from video clippings of the popular TV show, “Friends”. A seven way classification task is performed on this dataset, with each utterance being labeled as one of the $7$ emotions - “angry”, “sad”, “joy”, “neutral”, “fear”, “surprise” or “disgust”.
IV-B3 CMU-MOSI
The CMU-MOSI dataset [45] has a total of $2199$ utterances. Each utterance is labeled in the range of $[-3,3]$. Following previous works, we treat this as a binary classification problem with utterances having sentiment values in the range $[-3,0)$ being classified as negative sentiment and those with values in the range $[0,3]$ considered as positive sentiment. The dataset partitioning follows a prior work [51].
IV-B4 DAIC-WOZ
The DAIC-WOZ dataset [46] is a benchmark dataset for depression detection, consisting of $189$ clinical interviews between the patient and the interviewer. Out of these $189$ interviews, $107$ are part of the training set while $35$ interviews are part of the development subset. The dataset suffers from a data imbalance problem, with only $30$ interviews labeled as “depressed” in the train set . In order to increase the balance, we follow [52] and extract $100$ utterances randomly from each interview, which is labeled as depressed, while only $39$ utterances are selected for interviews classified as “normal”. The utterances from each interview are chosen randomly for the $5$ splits. Following prior work [53, 54, 55, 52], we report results on the development set of this dataset.
IV-B5 RAVDESS-Song
The RAVDESS-Song dataset [47] has a total of $1012$ song recordings by $23$ different singers. Each recording in this dataset is sung in one of six different emotions, namely, “neutral”, “calm”, “happy”, “sad”, “angry” and “fear”. We conduct a speaker independent evaluation for this dataset, and create $5$ different splits. For each split, we keep recordings from $16$ singers for training, while recordings from $3$ separate singers are used for validation. The recordings from the remaining $4$ speakers are used for evaluation.
IV-B6 CaFE
IV-B7 EmoDB
The EmoDB dataset [49] has a total of $535$ utterances spoken by 10 different speakers for the task of emotion recognition in German. Each utterance in this dataset is categorized as one of the seven emotions - “neutral”, “angry”, “disgust”, “sad”, “boredom”, “fear” and “joy”. Similar to RAVDESS-Song and CaFE, we create $5$ different speaker independent splits for this dataset. The utterances belonging to $6$ different speakers are chosen for training while the remaining $4$ speakers are used for validation and testing equally. The speakers used for train, validation and testing are chosen randomly for each of the $5$ splits.
IV-C Loss and Evaluation Metrics for Downstream Tasks
Since these datasets are categorical in nature, we use the cross entropy loss for training the downstream model weights (the convex combination weights and the lightweight classification head parameters). For testing, we use the weighted F1-score as the evaluation metric as many of the datasets are class-imbalanced in nature (See Table I) . Denoting the F1 score of a class $c$ with $N_{c}$ samples, by $F1_{c}$, the weighted F1-score is defined as:
$WF1=\frac{1}{\sum_{c=1}^{C}N_{c}}\sum_{c=1}^{C}N_{c}\times F1_{c}$ (3)
IV-D Implementation Details
IV-D1 Pre-training
During pre-training, all the speech utterances from MSP-PODCAST are padded or randomly cropped to a duration of $5$ seconds. The model is trained with a learning rate of $1e$-$5$ and a batch size of $128$ with AdamW [56] as the optimizer. The model is trained for a total of $200,000$ steps and the best model parameters based on validation set performance are chosen for evaluation of downstream datasets.
IV-D2 Fine-tuning and evaluation
For the downstream task training, the speech signals are cropped to a maximum duration of $30$ seconds or padded to a minimum duration of $1$ second. For the depression detection dataset, DAIC-WOZ, each speech segment has a duration of $10$ seconds [52].
Each layer output in the common, semantic, and acoustic encoders has a dimensionality of $T\times 768$, where $T$ denotes the number of frames in the speech signal, sampled at $50$Hz. For the CARE model, as outputs from the $6$ semantic and acoustic encoder layers are concatenated, the combined output dimension is $6\times T\times 1536$. To align with this dimensionality, the output from the convolutional feature extractor and the common encoder’s $6$ layers are duplicated to yield features of dimension $7\times T\times 1536$. Representations from these $13$ layers are combined through a convex combination approach with learnable weights producing features of dimension $T\times 1536$. Following this, features are mean-pooled along the temporal dimension, producing a single $1536$-dimensional vector per audio file. This is input into a classification head consisting of a two-layer feed-forward neural network that employs ReLU activation [57]. Only the weights for the convex combination of layer representations and those in the two-layer fully connected classification head are trained on each downstream dataset, consistent with the SUPERB framework [41].
We use a batch size of $32$ with a learning rate of $1e$-$4$ and train the model for $50$ epochs. The hidden dimension of the two-layer classification head is set to be $256$. The AdamW optimizer is used here as well. All the models, including the CARE and the baseline systems, utilize the same classification backend. Thus, the design allows fair comparison of the different representations.  [^4] Code available at https://github.com/iiscleap/CARE.
TABLE II: Comparison with other works for downstream datasets. # indicates the models where the downstream dataset is contained in the pre-training dataset. Results in bold indicate the best performing model, while those underlined indicate the second-best model. All numbers are weighted F1-scores computed over $5$ random initializations (mean and standard deviation shown). Number of parameters used during inference are also mentioned. Vesper [20] is not compared due to unavailability of the model weights
Params:$94$M Params:$94$M Params:$94$M Params:$94$M Params:$600$M Params:$7$B Params:$13$B Params:$160$M
IEMOCAP-4 $65.85^{\pm 0.45}$ $64.96^{\pm 0.21}$ $62.68^{\pm 0.72}$ $67.45^{\pm 0.59\#}$ $59.43^{\pm 0.36}$ $\mathbf{75.78}^{\pm 0.55\#}$ $\underline{72.92}^{\pm 2.33\#}$ $69.39^{\pm 0.46}$
IEMOCAP-6 $51.67^{\pm 0.51}$ $50.72^{\pm 0.85}$ $46.01^{\pm 0.35}$ $54.05^{\pm 0.58\#}$ $43.48^{\pm 0.18}$ $\mathbf{59.31}^{\pm 1.39\#}$ $\underline{58.07}^{\pm 1.62\#}$ $55.02^{\pm 0.39}$
MELD $45.61^{\pm 0.36}$ $45.30^{\pm 0.55}$ $41.89^{\pm 0.52}$ $47.60^{\pm 0.33\#}$ $43.16^{\pm 0.19}$ $\mathbf{53.34}^{\pm 0.69}$ $\underline{52.59}^{\pm 0.39}$ $48.05^{\pm 0.79}$
CMU-MOSI $64.09^{\pm 0.77}$ $62.53^{\pm 0.55}$ $59.67^{\pm 0.35}$ $66.45^{\pm 0.62}$ $\underline{74.61}^{\pm 0.26}$ $\mathbf{78.03}^{\pm 0.72}$ $72.80^{\pm 1.03}$ $66.74^{\pm 0.99}$
DAIC-WOZ $63.15^{\pm 1.53}$ $65.87^{\pm 1.98}$ $\underline{67.81}^{\pm 1.37}$ $61.62^{\pm 0.65}$ $64.32^{\pm 0.37}$ $62.61^{\pm 3.42}$ $64.72^{\pm 3.02}$ $\mathbf{68.49}^{\pm 2.11}$
RAVDESS-Song $50.45^{\pm 3.56}$ $\underline{53.46}^{\pm 1.14}$ $38.53^{\pm 5.23}$ $48.49^{\pm 1.01}$ $11.81^{\pm 1.95}$ $50.20^{\pm 1.34}$ $51.92^{\pm 3.59}$ $\mathbf{60.11}^{\pm 1.57}$
CaFE $66.58^{\pm 2.55}$ $66.53^{\pm 4.51}$ $48.75^{\pm 4.32}$ $59.27^{\pm 3.78}$ $5.73^{\pm 1.35}$ $59.86^{\pm 2.02}$ $\underline{69.84}^{\pm 3.25}$ $\mathbf{76.98}^{\pm 1.45}$
EmoDB $66.54^{\pm 4.79}$ $66.94^{\pm 3.91}$ $48.87^{\pm 3.14}$ $64.37^{\pm 2.62}$ $10.2^{\pm 2.02}$ $\underline{82.80}^{\pm 2.93}$ $82.15^{\pm 3.96}$ $\mathbf{83.41}^{\pm 1.98}$
IV-E Performance of CARE
The results on the $8$ downstream datasets using representations from the proposed CARE model are shown in Table II. A comparative analysis with several recent studies is also presented. These baseline models are categorized into two groups based on the number of parameters used during inference: base models (parameter size $<200$M), and large models ($>500$M), which also include LLM based models. The following observations are made for each category:
IV-E1 Base models
We compare HuBERT [15], WavLM [32], data2vec [58] and emotion2vec [21] representations as the baseline models in this category. Among these baseline systems, the emotion2vec is also pre-trained on IEMOCAP and MELD datasets, partially explaining the improved results seen on the downstream tasks on these datasets. While CARE performs similar to emotion2vec on CMU-MOSI, it improves over all the base-sized models on other datasets. On the average, the proposed CARE achieves a relative improvement of $15.6$% over the best baseline model (HuBERT).
IV-E2 Large models
IV-E3 LLM based models
Two versions of SALMONN [30] ($7$B and $13$B) [^8] https://huggingface.co/tsinghua-ee/SALMONN are considered as examples of LLM-based models. These are typically applied in a zero-shot setting; however, due to variability in emotion classes across datasets, their zero-shot performance is inconsistent. E.g. while SALMONN-$13$B model achieves $68.75\%$ weighted F1-score on the IEMOCAP-4 dataset (on which it is trained), it achieves only $24.06\%$ for MELD. Thus, for fair comparison, the same framework used in CARE and other baseline models is followed for the LLM based evaluations as well. The internal representations from all layers ($41$ layers for SALMONN $13$B and $33$ layers for SALMONN $7$B) are aggregated using a convex combination, and the classification head (similar to CARE) is trained for each downstream dataset. Similar to emotion2vec, SALMONN includes IEMOCAP in its pre-training, leading to superior performance on IEMOCAP-4 and IEMOCAP-6 compared to CARE. The larger model size and extensive pre-training data allows SALMONN to outperform CARE by $10\%$ and $34\%$ (relative improvements) on the MELD and CMU-MOSI datasets, respectively. However, on the remaining four tasks, CARE surpasses the SALMONN models, achieving relative improvements of $17\%$ and $24\%$ on the RAVDESS-song and CaFE datasets, respectively. Notably, though music datasets are used to pre-train SALMONN, CARE emerges as the best model on the RAVDESS-Song dataset.
Key takeaways: 1) On average, CARE emerges as the top-performing model across the eight datasets, surpassing even the SALMONN $13$B model, which has nearly $80$ times more parameters. Although LLM-based models show strengths in in-domain emotion recognition datasets, their performance declines on out-of-domain tasks, indicating limited generalizability across diverse tasks and multilingual emotional speech. 2) CARE’s advantage over speech SSL models like WavLM, HuBERT, and data2vec is expected, given that these models are trained on non-emotional data (see Sec. V-F for a related experiment). 3) Notably, CARE outperforms the multilingual SONAR model on CaFE and EmoDB datasets although it is trained on English speech only. This showcases the generalizability of our pre-training technique to out-of-domain tasks in SER.
IV-F Emotional Attribute Prediction
The emotion recognition can be posed as a regression problem, where valence, arousal and dominance of a particular utterance are predicted [59]. We use the MSP-IMPROV [60] for this purpose. This is an audio-visual dataset that consists of $12$ actors eliciting a set of sentences in different emotions. The dataset consists of $8438$ utterances with valence, arousal and dominance values (ranging from $1$ to $5$). We split the dataset in $12$ parts, where each part contains utterances corresponding to $10$ training speakers, while speech from the two other speakers are used for validating and testing the model. The performance is measured as the average over these $12$ parts.
In Eq. 4, $\rho$ is the Pearson’s correlation coefficient between the ground truth and the predicted scores. For training the downstream model, the representations from CARE and other models are aggregated similar to the categorical datasets. This is followed by a two-layer regression head with $256$ as the hidden dimension and $3$ as the output dimension ($1$ for each of the three attributes). The objective is to increase the CCC between the ground truth and the predicted values for each of the dimensions of valence, arousal and dominance. In other words, the loss used for each attribute is $L_{ccc}=1-CCC$. The results for this dataset along with other baseline models are shown in Table.III. We note that for this task, the CARE embeddings achieve the best results in terms of the valence and dominance attributes, while the performance on arousal is marginally better for the SALMONN-$7$B model.
TABLE III: Results for the MSP-IMPROV dataset. CCC stands for Concordance Correlation Coefficient while V, A, D stand for valence, arousal and dominance respectively.
Method CCC-V CCC-A CCC-D
WavLM-base [16] $0.51$ $0.64$ $0.47$
emotion2vec [21] $0.5$ $0.61$ $0.49$
SALMONN-$7$B [30] $0.53$ $\mathbf{0.67}$ $\underline{0.52}$
SALMONN-$13$B [30] $\underline{0.56}$ $0.65$ $0.51$
CARE $\mathbf{0.57}$ $\underline{0.66}$ $\mathbf{0.53}$
V Discussion
@@
TABLE IV: Baseline results on the different downstream datasets in terms of weighted F1-score. All numbers are averaged over $5$ random initializations of the downstream network. We also show the results of the different components of CARE in this table. Number of parameters used during inference are also mentioned
@@
Semantic +
Common Enc.
Acoustic +
Common Enc.
CARE
@@
V-A Comparison with Baselines
The performance of CARE is further compared with two baselines based on the supervisory signals used during pre-training. The results are shown in Table IV.